OpenStudy (anonymous):

i forgot there is 2 more parts to this question!! :( A manufacturer of lawn mowers uses the function A(n)=(132n+75250)/n to model the average cost per lawn mower, in dollars, where n is the number of lawn mowers produced. b. Describe the meaning of the horizontal asymptote in the context of the problem. c. What is the minimum number of lawn mowers that must be manufactured to bring the average cost per lawn mower down to $199? Round appropriately and state the answer in a complete sentence.

jimthompson5910 (jim_thompson5910):

what is the horizontal asymptote in this case?

OpenStudy (anonymous):

132?

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

what this means is that as n gets really really large, A(n) will slowly start to approach 132

jimthompson5910 (jim_thompson5910):

so if you produce a very large amount of mowers, then the average cost will start to drift towards $132 and the most important thing is that the cost will never dip below $132 because it never crosses the horizontal asymptote in this case

jimthompson5910 (jim_thompson5910):

so that's like the wall that you approach, but never touch or cross

jimthompson5910 (jim_thompson5910):

make sense?

OpenStudy (anonymous):

yes... it does!:)

jimthompson5910 (jim_thompson5910):

ok great

jimthompson5910 (jim_thompson5910):

the next part will have you plug in A(n) = 199 and solving for n

jimthompson5910 (jim_thompson5910):

A(n)=(132n+75250)/n 199=(132n+75250)/n 199n=132n+75250 ... ... ... n = ??

OpenStudy (anonymous):

ok ill do it

jimthompson5910 (jim_thompson5910):

ok

OpenStudy (anonymous):

ugh, i think it is n=1123...if rounding up

OpenStudy (anonymous):

tank you so much

jimthompson5910 (jim_thompson5910):

I'm getting n = 1,123.1343283582 which rounds up to 1124

jimthompson5910 (jim_thompson5910):

you round up to guarantee you pass the 199 mark

OpenStudy (anonymous):

Oh yes, i thought under was better, but makes sense..thank you!!

jimthompson5910 (jim_thompson5910):

normally just rounding 1,123.1343283582 would give you 1123, but it would make A(n) less than 199 so that's why you have to round up

jimthompson5910 (jim_thompson5910):

yw